A company has a fleet of cars available for rent. Statistics that the Mercedes AMG G63 is the most rented car among their fleet and one in four customers request this make


If 5 of todays rentals are selected, what ist the probability that 2 clients would have rented AMG G63?


1
Expert's answer
2021-05-04T12:33:15-0400

"\text{Let A event that is rented Mercedes AMG G63}"

"\text {the probability of the occurrence of the event A:}"

"p(A) =\frac{1}{4} =0.25;p =0.25"

"\text{probability that the event A will not occur in this test:}"

"\bar{p(A)}= q(A) = 1-p = 0.75"

"\text{Select 5 rental options for today is an independent test}"

"\text{apply the Bernoulli formula:}"

"p{2 \atop 2}(A)=C{2 \atop 2}p^2q^3= 10*0.25^2*0.75^3\approx 0.26"

Answer:0.26